MARTHA MICRO EXAM 9 pd-Adobe Aciobat Reader DC Fue Edn View Wadow Help MARTHA MICRO E jamy Home Tools Question 2. Monopoly (4 points) Use the following to answer questions below Output Price Marginal Costs 56 1 $10 2 $9 5 $%6 $4 8 $3 $6 $6 $6 $6 $6 $6 A) What is the profit maximizing price and quantity for this monopolist? (2 B) If the marginal cost would have been $O, what would the profit maximizi quantity be? (2 P)
Answer #1

Solution:

A) Profit maximising condition for monopolist is MR = MC (Marginal revenue = Marginal cost)

At quantity = 3, MR = (3*8)-(2*9) = $6 and MC = $6

Hence, profit maximising quantity = 3 and corresponding price = $8

B) If marginal cost would have 0 then the monopolist would produce 6 units because at this point MR = (6*5)-(5*6) = 0 =MC.